OpenStudy (anonymous):

(PLEASE HELP, I get confused in these types of questions and end up getting them wrong, please help) A music store offers two versions of a popular song. the size of the standard version is 2.4 MB. the size of the High-quality version is 4.2 MB. Yesterday there were 1240 downloads of the song, for a total download size of 3678 MB. How many downloads of the standard version were there?

OpenStudy (anonymous):

2.4(x) + 4.2(1240-x) = 3678

OpenStudy (anonymous):

x is the number of low quality downloads, 1240-x is the number of high quality downloads. you are multiplying each quantity by their respective size, and summing to the total download size of 3678 mb

OpenStudy (anonymous):

After distributing, your equation is: \[2.4x + 5208 - 4.2x = 3678\]

OpenStudy (anonymous):

Combining like terms: \[5208-1.8x = 3678\] Isolating x: \[1.8x = 5208-3678\]\[1.8x = 1530\]\[x = \frac{1530}{1.8}\]\[x = 850\] So, there were 850 low quality downloads.

OpenStudy (anonymous):

S = standard Q = high-quality S+Q = 1240 total downloads 2.4 S + 4.2 Q = 3678 total Mbytes Find S and Q. Eliminate S temporarily by multiplying the first equation by -2.4 and adding the two equations together to get an equation in Q alone -2.4 S -2.4 Q = (-2.4)(1240) = -2976 2.4 S + 4.2 Q = 3678 total Mbytes gives 1.8 Q = 702 Q = 702/1.8 = 390 downloads S = 1240 - 390 = 850 downloads There you are.

OpenStudy (anonymous):

Since there are 1240-x high quality downloads, x = 850, meaning there are 1240-850 high quality download = 390 high quality downloads.

OpenStudy (anonymous):

Thank you both so much.

OpenStudy (anonymous):

You're welcome :) Do you understand how to set up these problems now?

OpenStudy (anonymous):

Yes, you both gave me a better understanding of it. Thanks again.

OpenStudy (anonymous):

Great, have a nice day!!